OpenStudy (geneticrockhopper247):

Okay, so I'm having a little trouble understanding understanding these directions for a Physics project: Calculate the average acceleration. Find the change in speed between the first and second time points and the second and third time points and divide this by the change in time. You may want to add another row to your data tables for these calculations. Can someone please help me???

OpenStudy (geneticrockhopper247):

I NEED THIS BY TONIGHT!!!!

OpenStudy (anonymous):

use your basic kinematic equations: v1=vo+at v1^2=vo^2+2ad x=vo*t-0.5*a*(t^2)

OpenStudy (anonymous):

it looks like your teacher wants you to find each velocity between each point interval. once you do this you have your change in velocity. Subtract your largest time interval (last point probably) from your initial time (first point probably). This gives you the change in time. once you have those values calculate the acceleration with \[\frac{ \Delta v}{ \Delta t} = a\]

OpenStudy (geneticrockhopper247):

Thank you guys so much!!!!! I really appreciate it!